OpenStudy (anonymous):

PLEASE HELP ME!!! I JUST NEED TO KNOW HOW TO DO THIS!! Verify the identity. cos(4u)=cos^2(2u) - sin^2(2u)

OpenStudy (anonymous):

I made this trig sheet. http://learnix.net/wordpress/wp-content/uploads/Trig-Cheat-Sheet-1.4.pdf Can you see an identity on it that might fit this situation?

OpenStudy (anonymous):

I'm really not sure how this is done. This is why I am asking how it can be done

OpenStudy (anonymous):

check out on my sheet where it says Angle Sum and Difference Do you see the cosine one, which is second from top?

OpenStudy (anonymous):

yes? Could you please tell me how to do it?

OpenStudy (anonymous):

like what steps would I take?

OpenStudy (anonymous):

So what do we need to do to make cos(4u) to look like cos(x+y) where x and y are anything. Sorry for the delayed response too.

OpenStudy (anonymous):

i don't know, please just explain to me

OpenStudy (anonymous):

okay, so we can rewrite cos(4u) as cos(2u+2u) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

According to my cheat sheet you follow the right hand side of the equation now. Can you do that?

OpenStudy (anonymous):

no, please just show me how to do it. I'm a visual learning

OpenStudy (anonymous):

@Hero @geny55 @deana99 @omarbirjas @johnweldon1993 @Awolflover1 @Quan99 @EclipsedStar @razor99 @flvsucks @Nnesha @clamin @xo.minnie.xox @zepdrix @Pocarii @pixy @highjumpcollect @water_wings @badmood @BellaWolfe @Yolo4mecuite @yisrael012332 @toxicsugar22 @ozhobbits @graysongraddylol

OpenStudy (loser66):

\(cos (4u) = cos (2u+2u)\) ok?

OpenStudy (anonymous):

okay, why?

OpenStudy (loser66):

not yet, be patient

OpenStudy (anonymous):

you got it

OpenStudy (loser66):

cos (a+b) = cos acosb - sina sinb now, apply to your problem with a = 2u and b = 2u also you have : \(cos (4u) = cos (2u+2u) = cos (2u)cos (2u) - sin(2u) sin(2u) = cos^2 (2u) -sin^2 (2u)\) dat sit.

OpenStudy (anonymous):

Thank you!!